### v1

@@ 5. The Stokes operator
    For each $n\in\mathrm{N}$, $(\mathbf{v}_{n},\omega_{n})$ is an eigenpair of the Stokes operator, i.e.,
    $A\mathbf{v}_{n}=\omega_{n}\mathbf{v}_{n},\qquad\text{on $\Gamma$}.$ (5.6)
@@
    $\{\omega_{n}^{-1/2}\mathbf{v}_{n}\}_{n\in\mathrm{N}}$ is an orthonormal basis of $\mathbf{H}^{1}_{t,\sigma}(\Gamma)$ with respect to the equivalent inner product $(\mathbf{u},\mathbf{v})\mapsto\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v}$ (cf. 3.8 Poincaré inequality).
  4. (iv)
@@
Due to (5.1), 2.10 and 5.2 (properties of the Stokes operator), the operator $A^{-1}$ from $\mathbf{L}^{2}_{t}(\Gamma)$ to $\mathbf{L}^{2}_{t}(\Gamma)$ is compact, symmetric and positive. Hence, since $\mathbf{L}^{2}_{t}(\Gamma)$ is separable, it follows from [28, §D.6] that there exist an orthonormal basis $\{\mathbf{v}_{n}\}_{n\in\mathrm{N}}$ of $\mathbf{L}^{2}_{t,\sigma}(\Gamma)$ and a sequence of nonincreasing numbers $\{\lambda_{n}\}_{n\in\mathrm{N}}$ with $\lim_{n\to\infty}\lambda_{n}=0$ such that $A^{-1}\mathbf{v}_{n}=\lambda_{n}\mathbf{v}_{n}$, or equivalently, $A\mathbf{v}_{n}=\omega_{n}\mathbf{v}_{n}$ with $\omega_{n}=\lambda_{n}^{-1}$. For each $2<p<\infty$, the $\mathbf{W}^{m+2,p}_{t,\sigma}(\Gamma)$-regularity of each $\mathbf{v}_{n}$ follows from a bootstrapping argument resulting from combining the $\mathrm{L}^{p}$-based Sobolev regularity of the tangent Stokes equations (cf. 4.9) with Sobolev embeddings (cf. Section 2.2) and iterating as many times as the regularity of $\Gamma$ permits. Finally, the $\mathbf{H}^{1}_{t,\sigma}(\Gamma)$-orthonormality of $\{\omega_{n}^{-1/2}\mathbf{v}_{n}\}_{n\in\mathrm{N}}$ follows from 5.2; namely
$\int_{\Gamma}\nabla_{\Gamma}\mathbf{v}_{n}:\nabla_{\Gamma}\mathbf{v}_{k}=\int_{\Gamma}A\mathbf{v}_{n}\cdot\mathbf{v}_{k}=\omega_{n}\int_{\Gamma}\mathbf{v}_{n}\cdot\mathbf{v}_{k}=\omega_{n}\delta_{nk},\qquad\forall n,k\in\mathrm{N}.$ (5.7)
To show that $\{\mathbf{v}_{k}\}_{k\in\mathrm{N}}$ is basis of $\mathbf{H}^{1}_{t,\sigma}(\Gamma)$, let $\mathbf{u}\in\mathbf{H}^{1}_{t,\sigma}(\Gamma)$ such that $\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v}_{k}=0$ for each $k\in\mathrm{N}$. Then, by (5.5) and (5.6), $\int_{\Gamma}\mathbf{u}\cdot A\mathbf{v}_{k}=\omega_{k}\int_{\Gamma}\mathbf{u}\cdot\mathbf{v}_{k}=0$ for each $k\in\mathrm{N}$. Since $\omega_{k}\neq 0$, we deduce by (missing) i that $\mathbf{u}=\mathbf{0}$. Hence, $\operatorname{span}\{\omega_{n}^{-1/2}\mathbf{v}_{n}\}_{n\in\mathrm{N}}$ is dense in $\mathbf{H}^{1}_{t,\sigma}(\Gamma)$.
The proof of (missing) iv follows similarly to (missing) iii, by also using the bijectivity of $A$ and (5.7) if $m+2$ is odd. We omit further details. ∎
We can also define fractional powers of the Stokes operator $A$ as usual [19, p. 33].
Definition 5.4 (spectral Fractional Stokes operator on $\Gamma$).
Let $\alpha>0$. We define the operator $A^{\alpha}:D(A^{\alpha})\rightarrow\mathbf{L}^{2}_{t,\sigma}(\Gamma)$ by
$A^{\alpha}\mathbf{v}:=\sum_{n=1}^{\infty}\omega_{n}^{\alpha}\beta_{n}\mathbf{v}_{n},\qquad\forall\mathbf{v}=\sum_{n=1}^{\infty}\beta_{n}\mathbf{v}_{n}\in D(A^{\alpha}),$
where $D(A^{\alpha})$ is the linear subspace of $\mathbf{L}^{2}_{t,\sigma}(\Gamma)$ defined by
$D(A^{\alpha}):=\left\{\sum_{n=1}^{\infty}\beta_{n}\mathbf{v}_{n}\in\mathbf{L}^{2}_{t,\sigma}(\Gamma):\{\beta_{n}\}_{n\in\mathrm{N}}\subseteq\mathrm{R},\sum_{n=1}^{\infty}\omega_{n}^{2\alpha}\beta_{n}^{2}<\infty\right\}.$
We endow $D(A^{\alpha})$ with the inner product
$\langle\mathbf{v}\in D(A^{\alpha}),\mathbf{w}\rangle_{\alpha}:=\sum_{n=1}^{\infty}\omega_{n}^{2\alpha}\beta_{n}\gamma_{n},\qquad\forall\mathbf{v}=\sum_{n=1}^{\infty}\beta_{n}\mathbf{v}_{n},\quad\mathbf{w}=\sum_{n=1}^{\infty}\gamma_{n}\mathbf{v}_{n}\in D(A^{\alpha}).$ (5.8)
Remark 5.5.
For each $\alpha_{2}>\alpha_{1}>0$, the following embeddings are continuous
$D(A^{\alpha_{2}})\hookrightarrow D(A^{\alpha_{1}})\hookrightarrow\mathbf{L}^{2}_{t,\sigma}(\Gamma),$
because of the Parseval identity in $\mathbf{L}^{2}_{t,\sigma}(\Gamma)$, and $(\omega_{n}/\omega_{1})^{\alpha_{2}-\alpha_{1}}>1$ and $(\omega_{n}/\omega_{1})^{\alpha_{1}}>1$ for each $n\in\mathrm{N}$.
Proposition 5.6 (basis of $D(A^{\alpha})$).
Let $\alpha>0$. Then, the family of functions $\{\omega_{n}^{-\alpha}\mathbf{v}_{n}\}_{n\in\mathrm{N}}$ is an orthonormal basis of $D(A^{\alpha})$.
Proof.
The orthonormality of the family is easy to see. Now, let $\mathbf{v}=\sum_{n=1}^{\infty}\beta_{n}\mathbf{v}_{n}\in D(A^{\alpha})$ such that $\langle\mathbf{v},\omega_{m}^{-\alpha}\mathbf{v}_{m}\rangle_{\alpha}=0$ for each $m\in\mathrm{N}$. That is, by (5.8), $\omega_{m}^{2\alpha}\beta_{m}\omega_{m}^{-\alpha}=0$ for each $m\in\mathrm{N}$. Since $\omega_{m}\neq 0$, we deduce that $\beta_{m}=0$ for each $m$, and hence $\mathbf{v}=\mathbf{0}$. This proves that the orthogonal complement of $\operatorname{span}\{\omega_{n}^{-\alpha}\mathbf{v}_{n}\}_{n\in\mathrm{N}}$ $D(A^{\alpha})$ is $\{\mathbf{0}\}$. Hence, $\operatorname{span}\{\omega_{n}^{-\alpha}\mathbf{v}_{n}\}_{n\in\mathrm{N}}$ is dense in $D(A^{\alpha})$. ∎
The following result states that the fractional Stokes operator generalizes the multiple composition of the Stokes operator $A$ (cf. 5.1). Due to the absence of boundary conditions (because $\Gamma$ is closed) the statement is considerably simpler than, e.g., in [19, Proposition 4.12].
Proposition 5.7 (characterization of the fractional Stokes operator).
Let $\Gamma$ be of class $C^{\max\{3,m\}}$ for some integer $m\geq 1$. Then,
  1. (i)
    $D(A^{\frac{m}{2}})=\mathbf{H}^{m}_{t,\sigma}(\Gamma)$, and for each $\mathbf{v}\in D(A^{\frac{m}{2}})$ its defining series also converges in the $\mathbf{H}^{m}(\Gamma)$-norm.
  2. (ii)
    If $m$ is even, the fractional Stokes operator $A^{\frac{m}{2}}$ (cf. 5.4) coincides with the $\frac{m}{2}$-th power of the Stokes operator $A:\mathbf{H}^{2}_{t,\sigma}(\Gamma)\rightarrow\mathbf{L}^{2}_{t,\sigma}(\Gamma)$ (cf. 5.1).
  3. (iii)
@@
    For each $\mathbf{v}=\sum_{n=1}^{\infty}\beta_{n}\mathbf{v}_{n},\mathbf{w}=\sum_{n=1}^{\infty}\gamma_{n}\mathbf{v}_{n}\in\mathbf{H}^{m}_{t,\sigma}(\Gamma)$,
    $\langle\mathbf{v},\mathbf{w}\rangle_{m}=\begin{cases}\int_{\Gamma}A^{\frac{m}{2}}\mathbf{v}\cdot A^{\frac{m}{2}}\mathbf{w},&\text{if $m$ is even},\\ \int_{\Gamma}\nabla_{\Gamma}A^{\frac{m-1}{2}}\mathbf{v}:\nabla_{\Gamma}A^{\frac{m-1}{2}}\mathbf{w},&\text{if $m$ is odd}.\end{cases}$
Proof.
Let $\mathbf{v}=\sum_{n=1}^{\infty}\beta_{n}\mathbf{v}_{n}$ be an arbitrary element of $\mathbf{L}^{2}_{t,\sigma}(\Gamma)$. If $m\geq 0$ is even, we have the following set of equivalences,
[…]
where the last equivalence is due to [12, Lemma 5.1] and (missing) iv of 5.3 (eigendecomposition of the Stokes operator). Similarly, if $m\geq 1$ is odd, then
[…]
This proves (missing) i. The proof of (missing) ii follows easily from (5.6) and the continuity of $\underbrace{A\cdots A}_{\text{$\frac{m}{2}$ times}}$ on $\mathbf{H}^{m}_{t,\sigma}(\Gamma)$. Similarly, (missing) iii follows from (missing) ii and (5.7). We omit further details. ∎
Proposition 5.8 (algebraic properties).
Let $\alpha=\alpha_{1}+\alpha_{2}$ with $\alpha_{1},\alpha_{2}>0$. Then, for each $\mathbf{v},\mathbf{w}\in D(A^{\alpha})$,
$\int_{\Gamma}A^{\alpha_{1}}\mathbf{v}\cdot A^{\alpha_{2}}\mathbf{w}=\int_{\Gamma}A^{\alpha}\mathbf{v}\cdot\mathbf{w}=\int_{\Gamma}A^{\alpha_{2}}\mathbf{v}\cdot A^{\alpha_{1}}\mathbf{w},$ (5.9)
Proof.
Let $\mathbf{v}=\sum_{n=1}^{\infty}\beta_{n}\mathbf{v}_{n},\mathbf{w}=\sum_{n=1}^{\infty}\gamma_{n}\mathbf{v}_{n}\in D(A^{\alpha})\subseteq D(A^{\alpha_{1}})\cap D(A^{\alpha_{2}})$. Then,
$\int_{\Gamma}A^{\alpha_{1}}\mathbf{v}\cdot A^{\alpha_{2}}\mathbf{w}=\int_{\Gamma}\left(\sum_{n=1}^{\infty}\omega_{n}^{\alpha_{1}}\beta_{n}\mathbf{v}_{n}\right)\cdot\left(\sum_{n=1}^{\infty}\omega_{n}^{\alpha_{2}}\gamma_{n}\mathbf{v}_{n}\right)\\ =\sum_{n=1}^{\infty}\omega_{n}^{\alpha_{1}+\alpha_{2}}\beta_{n}\gamma_{n}=\int_{\Gamma}\left(\sum_{n=1}^{\infty}\omega_{n}^{\alpha}\beta_{n}\mathbf{v}_{n}\right)\cdot\left(\sum_{n=1}^{\infty}\gamma_{n}\mathbf{v}_{n}\right)=\int_{\Gamma}A^{\alpha}\mathbf{v}\cdot\mathbf{w}.$
Since the role of $\alpha_{1}$ and $\alpha_{2}$ is symmetric, this proves (5.9). ∎
@@ 6. Tangent Navier–Stokes equations
Given $1<p<\infty$ (to be determined later) and denoting $p^{*}:=\frac{p}{p-1}$ its dual exponent , we consider the following variational formulation associated to (NS): Given $\mathbf{f}\in(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}$ and $g\in\mathrm{L}^{p}_{\#}(\Gamma)$, find $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}_{\#}^{p}(\Gamma)$ such that
@@
The value of $p$ needs to be restricted so as to ensure that the integral $\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\cdot\mathbf{v}$ makes sense; namely, by Hölder’s inequality $\int_{\Gamma}|(\nabla_{\Gamma}\mathbf{u})\overline{\mathbf{u}}\cdot\mathbf{v}|\leq\|\nabla_{\Gamma}\mathbf{u}\|_{0,p;\Gamma}\|\mathbf{u}\,\mathbf{v}^{T}\|_{0,p^{*};\Gamma}$ this reduces to finding $1<p<\infty$ such that $\|\overline{\mathbf{u}}\,\mathbf{v}^{T}\|_{0,p^{*};\Gamma}$ is finite for each $\forall\overline{\mathbf{u}}\in\mathbf{W}^{1,p}_{t}(\Gamma)$ and $\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma)$. We argue first for $d=2$; Hölder’s inequality and Sobolev embeddings (cf. Section 2.2) imply
$\|\overline{\mathbf{u}}\,\mathbf{v}^{T}\|_{0,p^{*};\Gamma}$ $\leq\begin{cases}\|\overline{\mathbf{u}}\|_{0,4;\Gamma}\|\mathbf{v}\|_{0,4;\Gamma},&\text{if $p=2=p^{*}$},\\ \|\overline{\mathbf{u}}\|_{0,\Gamma}\|\mathbf{v}\|_{\mathbf{L}^{0,\frac{2p^{*}}{2-p^{*}}};\Gamma},&\text{if $p>2>p^{*}$},\\ \|\overline{\mathbf{u}}\|_{0,\frac{2p}{2-p};\Gamma}\|\mathbf{v}\|_{0,\frac{2p}{3p-4};\Gamma},&\text{if $\frac{4}{3}<p<2<p^{*}$},\\ \|\overline{\mathbf{u}}\|_{0,4;\Gamma}\|\mathbf{v}\|_{0,\infty;\Gamma},&\text{if $p=\frac{4}{3}$ (and so $p^{*}=4$)}\end{cases}$
$\leq C\|\overline{\mathbf{u}}\|_{1,p;\Gamma}\|\mathbf{v}\|_{1,p^{*};\Gamma},\qquad\forall\mathbf{u},\overline{\mathbf{u}}\in\mathbf{W}^{1,p}_{t}(\Gamma),\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma),$
as long as $p\geq\frac{4}{3}$.
For $d\geq 3$, the situation is a bit more complicated. If $d\leq p^{*}$ (which implies $d>p$), then it is easy to see that $\frac{dp}{d-p}<p^{*}$, and so Hölder’s inequality with Sobolev embeddings are not enough to control $\|\mathbf{u}\,\mathbf{v}^{T}\|_{0,p^{*};\Gamma}$. On the other hand, if $d>p^{*}$, then
$\|\mathbf{u}\,\mathbf{v}^{T}\|_{0,p^{*};\Gamma}\leq\|\overline{\mathbf{u}}\|_{0,d;\Gamma}\|\mathbf{v}\|_{0,\frac{dp^{*}}{d-p^{*}};\Gamma}\leq C\|\overline{\mathbf{u}}\|_{1,p;\Gamma}\|\mathbf{v}\|_{1,p^{*};\Gamma},\qquad\forall\mathbf{u},\overline{\mathbf{u}}\in\mathbf{W}^{1,p}_{t}(\Gamma),\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma)$
as long as $p\geq\frac{d}{2}$ (in particular, $p\geq\frac{d}{2}$ implies $d>p^{*}$). Overall, if $\frac{4}{3}\leq p<\infty$ if $d=2$ or $\frac{d}{2}\leq p<\infty$ if $d\geq 3$, then
$\int_{\Gamma}|(\nabla_{\Gamma}\mathbf{u})\overline{\mathbf{u}}\cdot\mathbf{v}|\leq\|\nabla_{\Gamma}\mathbf{u}\|_{0,p;\Gamma}\|\mathbf{u}\,\mathbf{v}^{T}\|_{0,p^{*};\Gamma}\leq C_{\operatorname{conv}}\|\mathbf{u}\|_{1,p;\Gamma}\|\overline{\mathbf{u}}\|_{1,p;\Gamma}\|\mathbf{v}\|_{1,p^{*};\Gamma},$ (6.2)

### v2

@@ 4.4. The Stokes operator
    For each $n\in\mathrm{N}$, $(\mathbf{v}_{n},\omega_{n})$ is an eigenpair of $A$:
    $A\mathbf{v}_{n}=\omega_{n}\mathbf{v}_{n},\qquad\text{a.e.~on $\Gamma$}.$
@@
Fix $p\in(2,\infty)$ and notice from 4.10 that each $\mathbf{v}_{n}\in\mathbf{W}^{1,p}_{t,\sigma}(\Gamma)$ satisfies (4.21), namely
$\int_{\Gamma}\nabla_{\Gamma}\mathbf{v}_{n}:\nabla_{\Gamma}\mathbf{w}=\omega_{n}\int_{\Gamma}\mathbf{v}_{n}\cdot\mathbf{w},\qquad\forall\mathbf{w}\in\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma).$
Hence, proceeding similarly to the proof of 4.5 (injectivity of the adjoint operator) we can utilize 4.3 (Helmholtz–Weyl decomposition on $\Gamma$) to show that there exists a unique $\uppi_{n}\in\mathrm{L}^{p}_{\#}(\Gamma)$ such that
$\int_{\Gamma}\nabla_{\Gamma}\mathbf{v}_{n}:\nabla_{\Gamma}\mathbf{w}-\int_{\Gamma}\uppi_{n}\operatorname{div}_{\Gamma}\mathbf{w}=\omega_{n}\int_{\Gamma}\mathbf{v}_{n}\cdot\mathbf{w},\qquad\forall\mathbf{w}\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma).$
Consequently, by iterating 4.8 (higher $\mathrm{L}^{p}$-based regularity of the tangent Stokes equations) as many times as the regularity of $\Gamma$ permits, we deduce that $\mathbf{v}_{n}\in\mathbf{W}^{m+2,p}_{t,\sigma}(\Gamma)$. Moreover, integration-by-parts formula (2.5) and properties of the Leray projection (cf. 4.4) yield $A\mathbf{v}_{n}:=-P_{\sigma}(\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}\mathbf{v}_{n})=\omega_{n}\mathbf{v}_{n}$ a.e. on $\Gamma$. This proves () i and () ii.
The proof of () iii follows similarly to the proof of () iv in 4.10, by also using the bijectivity of $A$ and the first equality of (4.23) if $m+2$ is odd. We omit further details. ∎
@@ 5. Tangent Navier–Stokes equations
In this section, we utilize the well-posedness, regularity and spectral theory for the (S) developed in Section 4, in conjunction with the Fredholm Alternative, to derive a corresponding well-posedness and regularity theory for the tangent Navier–Stokes equations (NS). More precisely, given $p\in[p_{0},\infty)$ ($p_{0}$ to be determined later), we are interested in the following variational formulation associated to (NS): Given $\mathbf{f}\in(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}$ and $g\in\mathrm{L}^{p}_{\#}(\Gamma)$, find $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}_{\#}^{p}(\Gamma)$ such that
@@
In contrast with (S), the presence of the nonlinear term $(\nabla_{\Gamma}\mathbf{u})\mathbf{u}$ in (NS) makes it necessary to restrict the range of $p$ to ensure that the integral $\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\cdot\mathbf{v}$ makes sense, as it also happens in the flat case; see, e.g., [9]. We now determine such range by first invoking Hölder’s inequality, yielding $\int_{\Gamma}|(\nabla_{\Gamma}\mathbf{u})\overline{\mathbf{u}}\cdot\mathbf{v}|\leq\|\nabla_{\Gamma}\mathbf{u}\|_{0,p;\Gamma}\|\overline{\mathbf{u}}\,\mathbf{v}^{T}\|_{0,p^{*};\Gamma}$. It remains to make sense of the latter for each $\overline{\mathbf{u}}\in\mathbf{W}^{1,p}_{t}(\Gamma)$ and $\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma)$. We argue first for $d=2$. Hölder’s inequality and Sobolev embeddings (cf. Section 2.1) imply
$\|\overline{\mathbf{u}}\,\mathbf{v}^{T}\|_{0,p^{*};\Gamma}$ $\leq\begin{cases}\|\overline{\mathbf{u}}\|_{0,4;\Gamma}\|\mathbf{v}\|_{0,4;\Gamma},&\text{if $p=2=p^{*}$},\\ \|\overline{\mathbf{u}}\|_{0,\Gamma}\|\mathbf{v}\|_{0,\frac{2p^{*}}{2-p^{*}};\Gamma},&\text{if $p>2>p^{*}$},\\ \|\overline{\mathbf{u}}\|_{0,\frac{2p}{2-p};\Gamma}\|\mathbf{v}\|_{0,\frac{2p}{3p-4};\Gamma},&\text{if $\frac{4}{3}<p<2<p^{*}$},\\ \|\overline{\mathbf{u}}\|_{0,4;\Gamma}\|\mathbf{v}\|_{0,\infty;\Gamma},&\text{if $p=\frac{4}{3}$ (and so $p^{*}=4$)}\end{cases}$
$\lesssim\|\overline{\mathbf{u}}\|_{1,p;\Gamma}\|\mathbf{v}\|_{1,p^{*};\Gamma},\qquad\forall\overline{\mathbf{u}}\in\mathbf{W}^{1,p}_{t}(\Gamma),\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma),$
as long as $p\geq\frac{4}{3}$. Conversely, the limiting bound $\|\overline{\mathbf{u}}\,\mathbf{v}^{T}\|_{0,4;\Gamma}\leq\|\overline{\mathbf{u}}\|_{0,4;\Gamma}\|\mathbf{v}\|_{0,\infty;\Gamma}$ for $p=\frac{4}{3}$ above makes it apparent that $\|\overline{\mathbf{u}}\,\mathbf{v}^{T}\|_{0,p^{*};\Gamma}$ cannot be controlled if $p<\frac{4}{3}$.
For $d\geq 3$, we proceed differently. If $d>p^{*}$, then Hölder’s inequality yields $\|\overline{\mathbf{u}}\,\mathbf{v}^{T}\|_{0,p^{*};\Gamma}\leq\|\overline{\mathbf{u}}\|_{0,d;\Gamma}\|\mathbf{v}\|_{0,\frac{dp^{*}}{d-p^{*}};\Gamma}$. If we further assume that $p\geq\frac{d}{2}$ (which actually implies $d>p^{*}$), then a Sobolev embedding lets us write
$\|\overline{\mathbf{u}}\,\mathbf{v}^{T}\|_{0,p^{*};\Gamma}\lesssim\|\overline{\mathbf{u}}\|_{1,p;\Gamma}\|\mathbf{v}\|_{1,p^{*};\Gamma},\qquad\forall\overline{\mathbf{u}}\in\mathbf{W}^{1,p}_{t}(\Gamma),\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma).$
On the other hand, if $3\leq d\leq p^{*}$ (which implies $d>p$), then it is easy to see that $\frac{dp}{d-p}<p^{*}$ and so it is impossible to exploit Hölder’s inequality in conjunction with the embedding $\mathbf{W}^{1,p}(\Gamma)\hookrightarrow\mathbf{L}^{\frac{dp}{d-p}}$ to control $\|\overline{\mathbf{u}}\,\mathbf{v}^{T}\|_{0,p^{*};\Gamma}$.
Consequently, we have deduced that if
$p\in\begin{cases}[\frac{4}{3},\infty)&\text{if $d=2$},\\ [\frac{d}{2},\infty)&\text{if $d\geq 3$},\end{cases}$ (5.2)
then
$\int_{\Gamma}|(\nabla_{\Gamma}\mathbf{u})\overline{\mathbf{u}}\cdot\mathbf{v}|\leq\|\nabla_{\Gamma}\mathbf{u}\|_{0,p;\Gamma}\|\overline{\mathbf{u}}\,\mathbf{v}^{T}\|_{0,p^{*};\Gamma}\leq C_{\operatorname{conv}}\|\mathbf{u}\|_{1,p;\Gamma}\|\overline{\mathbf{u}}\|_{1,p;\Gamma}\|\mathbf{v}\|_{1,p^{*};\Gamma},$ (5.3)